How many solutions does cosecant squared of x = 25 have? How do you find it?
Answer by stanbon(75887)

How many solutions does cosecant squared of x = 25 have? How do you find it?
------------------
If cscx = 25, sinx = 1/25 or 0.25
x = sin^-1 0.25 = 14.48 degrees or x = 180-14.48 = 165.52 degrees
--------
All the angles:
x = 14.48 + 360n where n= 0,1,2,3,,,,
x = 165.52 + 360n where n= 0,1,2,3,,,,
======================================
Cheers,
Stan H.